# EmMixformer: Mix transformer for eye movement recognition

### v1

@@ I-C Motivation and Contribution
Traditional approaches use manually-designed algorithms to extract, from eye movement, handcrafted features, input to ML approaches such as SVM and MLP, which suffers from the following issues: 1) Manually-designed models rely on human prior assumptions such as Gaussian distributions. These assumptions, however, are not always effective to learn robust eye movement patterns as the latter’s underlying data distributions are usually much more complex in real-life applications. This explains also why it is hard to design a mathematical model to effectively model such distributions; 2) Traditional ML approaches such as SVM and RF can be viewed as a neural network with very few (one or two) layers for feature representation learning. Their representation capacity is therefore relatively limited. DL models trained on huge data, by contrast, are capable of learning high-level features by stacking a number of layers. For eye movement recognition, DL models, such as CNN, have recently achieved promising performance compared to ML approaches and become an increasing modeling trend. CNN can effectively capture local patterns and features, such as edges, corners, textures, or shapes, by sharing weights and local perception. CNN, however, fail to capture temporal dependencies within sequences. As eye movement is typically temporal data, collected over time with observations recorded at different time points, existing DL-based approaches cannot extract effective features that embed such long-range time dependencies. Recently, transformers with attention mechanisms have emerged as new start-of-the art classifiers thanks to their high capacity to capture long dependencies among tokens, with impressive performance obtained, first, in various natural language processing (NLP) tasks [41] and then in computer vision [42].
Inspired by these facts, we propose in this paper a novel mix transformer for eye movement recognition, named EmMixformer. EmMixformer includes three modules, transformer, attention-LSTM, and Fourier-transformer. The eye movement signal is first split into fast and slow velocity signals, separately input to encoder for feature embedding. The embedding features are then input to the mix transformer module to learn a robust representation for classification. The transformer and attention LSTM aim at learning long and short temporal dependencies within the eye movement signal, while the Fourier transformer learns a global feature representation. Our mix transformer can, therefore, learn complementary features. The contributions of our work are summarized as follows:
@@ II-A Siamese CNN block
The Siamese CNN block consists of two sub-networks, each stacking four one-dimensional convolutional modules for feature extraction. Each module comprises a convolutional layer, a batch normalization layer, a ReLU layer, and an average pooling layer. The two subnets take respectively the preprocessed fast and slow eye movement signals as inputs and output the corresponding embedding feature representations. We then concatenate the output feature representations from the two subnets to obtain the complete feature representation.
@@ II-B Mix block
The Mix block comprises three modules, an attention LSTM (attLSTM), a Transformer, and a Fourier-transformer (Fourierformer). The three modules take the Siamese CNN block’s output as input to extract rich and comprehensive temporal-spatial features. Concretely, a temporal sequence from the Siamese CNN block is divided into a number of sub-sequences, which are input to attLSTM and transformer to extract the temporal features. The attLSTM module learns short-term dependencies within the temporal sequences, while the transformer module extracts long-term dependencies within the temporal sequences thanks to its self-attention mechanism. Their combination, therefore, can learn a comprehensive temporal feature representation. The third module, Fourierformer, takes as input the temporal sequences to learn a global frequency-domain feature representation. As a result, our mix block extracts both a feature representation embedding temporal dependencies and a global feature representation, thus effectively improving the modeling of time-series eye movement classification.
@@
Transformer, introduced in 2017 by Vaswani et al [41], has become the state of the art for NLP [41] and more recently computer vision [43]. Unlike traditional sequence models such as recurrent neural networks (RNNs) or CNNs, Transformer relies solely on a key idea, the self-attention mechanism, eliminating the need for sequential processing. The self-attention mechanism allows the model to weigh the importance of different tokens in a sequence, and enables it to capture effectively long-range dependencies and context information. By attending to all tokens simultaneously, Transformer can process sentences in parallel, leading to significantly faster training and inference processing times, compared to existing sequential models.
@@
where $f_{t}$ , $i_{t}$ and $o_{t}$ are the forgetting gate, input gate and output gate, respectively. $\widetilde{C_{t}}$ , $C_{t}$ and $h_{t-1}$ are the candidate cell state, the cell state, and the hidden state at the last moment, respectively. $\odot$ denotes the element-by-element matrix multiplication.
@@
$\mathcal{F}(X)\left(v\right)=\frac{1}{\sqrt{T}}\sum_{t=0}^{T-1}X(t)e^{-j2\pi\frac{vt}{T}}t$ (24)
where $\mathcal{F}(X)\left(v\right)$ is the complex value of the $v$-th frequency component in the frequency domain, $X(t)$ is the complex value of the $t$th sampling point in the time domain, $T$ is the sequence length, and $\frac{1}{\sqrt{T}}$ ensures normalized consistency between the DFT and the inverse Fourier transform (IDFT).
Using the real part $Re(X)$ and the imaginary part $Im(X)$ of $\mathcal{F}(X)\left(v\right)$, we obtain the amplitude component $A(X)$ and the phase component $\varphi\left(X\right)$ by the following equations:
$A(X)\left(v\right)=\sqrt{Re\left(X\right)\left(v\right)^{2}+Im\left(X\right)\left(v\right)^{2}}$ (25)
$\varphi\left(X\right)\left(v\right)=arctan\left(\frac{Im\left(X\right)\left(v\right)}{Re\left(X\right)\left(v\right)}\right)$ (26)
Next, we extract the feature representation from amplitude component $A(X)\left(v\right)$ and phase component $\varphi(X)\left(v\right)$ using two independent transformers. The self attention operation is performed to capture the amplitude information $A_{transformer}$ and phase information $\varphi_{transformer}$, computed as:
@@
Subsequently, the feature representations of the amplitude and phase components are combined into complex numbers and transformed back into the spatial domain by IFFT $\mathcal{F}^{-1}$:
@@ III-A Dataset
CEM-Ⅰ[23] Tobii TX300 300 yes 22
CEM-Ⅱ[23] EyeLink 1000 1000 yes 32
@@ III-D Experimental results for short time interval
These experiments aim to assess the verification performance of our approach on short time interval data. In the three datasets, the participants provided data from two sessions at each round. As detailed in section III-C, we select data from the first session for training and the remaining data for testing. The EMglasses dataset comprises 203 sequences in the training set and testing set, the four sub-datasets (RAN, TEX, FXS, and HSS) in the the GazeBase dataset comprises 322 sequences in the training set and the testing set, while the JuDo1000 dataset comprises 450 sequences in the training set and 150 sequences in the testing set. Note that in all datasets, we split each sequence, collected for several seconds, into several sub-sequences for training and testing. The state of the art approaches, i.e. DEL[38], Expansion CNN[7], Dense LSTM[36] and DenseNet [40] are employed for benchmarking. The EER and FRR@FAR of each approach are listed in Table III, Table IV, Table VII, Table V, Table VI, and Table VIII. The corresponding ROC curves are shown in Fig.7. It can be observed that our EmMixFormer outperforms the existing approaches and achieves the lowest verification error, i.e. 0.0655, 0.0672, 0.0795, 0.1660 on the four subsets of the GazeBase dataset, 0.0543 on the JuDo1000 dataset, and 0.1605 on our dataset. Also, we observe from Fig.7 that our proposed approach achieves the higher recognition accuracy w.r.t existing approaches at different FARs.
@@
EmMixFormer 0.0655 0.0234 0.1719 0.4062
DEL[38] 0.1333 0.1211 0.5391 0.8984
Expansion CNN[7] 0.1449 0.1289 0.5430 1.0000
Dense LSTM[36] 0.1078 0.0703 0.4688 0.7930
DenseNet[40] 0.1035 0.0625 0.4062 0.7812
@@
EmMixFormer 0.0672 0.1172 0.3516 0.5156
DEL[38] 0.1435 0.1680 0.7070 1.0000
Expansion CNN[7] 0.1271 0.1172 0.6953 0.9570
Dense LSTM[36] 0.0954 0.0703 0.4297 0.8164
DenseNet[40] 0.0886 0.0273 0.4414 0.8125
@@
EmMixFormer 0.0795 0.0312 0.3359 0.7109
DEL[38] 0.1565 0.2031 0.8008 0.9766
Expansion CNN[7] 0.1427 0.1739 0.7391 0.8594
Dense LSTM[36] 0.1669 0.2188 0.8281 0.9766
DenseNet[40] 0.1132 0.0859 0.5781 0.9023
@@
EmMixFormer 0.1660 0.1731 0.7891 0.9593
DEL[38] 0.2031 0.3242 0.8594 0.9844
Expansion CNN[7] 0.2165 0.4141 0.9258 1.0000
Dense LSTM[36] 0.2203 0.4805 0.9453 0.9883
DenseNet[40] 0.1886 0.2969 0.8218 0.9609
@@
EmMixFormer 0.1605 0.2156 0.7936 0.9603
DEL[38] 0.1955 0.3984 1.0000 1.0000
Expansion CNN[7] 0.2308 0.3594 0.9973 1.0000
Dense LSTM[36] 0.2332 0.5039 0.9805 0.9961
DenseNet[40] 0.1703 0.3301 0.8058 0.9612
@@ III-E Experimental results for long time intervals
TABLE IX: Results of long-term interval experiments
@@
EmMixFormer 0.0795 0.1802 0.2150 0.2133 0.2157 0.2717
DEL[38] 0.1565 0.2450 0.2807 0.3335 0.2765 0.3161
Expansion CNN[7] 0.1427 0.2947 0.3380 0.2895 0.3341 0.3728
Dense LSTM[36] 0.1690 0.2630 0.2880 0.3053 0.2633 0.3343
DenseNet[40] 0.1132 0.1957 0.2503 0.2406 0.2441 0.2913
@@ III-F Ablation experiment
The Mix block of our EmMixFormer consists of three modules: Attention LSTM, Transformer, and Fourier Transformer. To investigate the effect of each module w.r.t recognition accuracy improvement, we conducted ablation experiments on the round 1 data from the RAN sub-dataset in GazeBase. To ease the description, we remove attention LSTM and Fourier Transformer blocks from Mix block (as shown in Fig.2) and denote the resulting model as Transformer. Then, we gradually incorporate LSTM and Attention LSTM, with the resulting models denoted as LSTM+Transformer and attLSTM+Transformer, respectively. Afterwards, the Fourier Transformer, is included to the attLSTM+Transformer, with the resulting model represented as attLSTM+Transformer+Fourierformer, i.e. EmMixFormer. The verification error rates of these ablation schemes are shown in Table X and Fig.8. The experimental results demonstrate that the combination of the Attention LSTM and the Fourier Transformer allows to improve the identification accuracy of simple Transformer. Besides, comparing Transformer with LSTM+Transformer, we observe that the LSTM can provide compensation features, i.e. short dependencies to improve the performance of Transformer, which is supported by the work [53]. From Table X and Fig.8, we also observe that attLSTM+Transformer achieves lower EER than LSTM+Transformer, which implies that the attention mechanism improves LSTM performance. Finally, the performance becomes optimal when attLSTM+Transformer is combined with Fourier Transformer, which demonstrates that the attention mechanism in the frequency domain captures global dependencies for feature representation.
TABLE X: Results of the ablation experiments
Model Operation EER
EmMixFormer Transformer 0.1220
LSTM+Transformer 0.1023
attLSTM+Transformer 0.0897
attLSTM+Transformer+Fourierformer 0.0795

### v2

@@ I-C Motivation and Contribution
Traditional approaches use manually-designed algorithms to extract, from eye movement, handcrafted features, input to ML approaches such as SVM and MLP, which suffers from the following issues: 1) Manually-designed models rely on human prior assumptions such as Gaussian distributions. These assumptions, however, are not always effective to learn robust eye movement patterns as the latter’s underlying data distributions are usually much more complex in real-life applications. This explains also why it is hard to design a mathematical model to effectively model such distributions; 2) Traditional ML approaches such as SVM and RF can be viewed as a neural network with very few (one or two) layers for feature representation learning. Their representation capacity is therefore relatively limited. DL models trained on huge data, by contrast, are capable of learning high-level features by stacking a number of layers. For eye movement recognition, DL models, such as CNN, have recently achieved promising performance compared to ML approaches and become an increasing modeling trend. CNN can effectively capture local patterns and features, such as edges, corners, textures, or shapes, by sharing weights and local perception. CNN, however, is difficult to capture long temporal dependencies within sequences because of limited receptive field[41][42]. As eye movement is typically temporal data, collected over time with observations recorded at different time points, existing DL-based approaches is difficult to extract effective features that embed such long-range time dependencies. Recently, transformers with attention mechanisms have emerged as new start-of-the art classifiers thanks to their high capacity to capture long dependencies among tokens, with impressive performance obtained, first, in various natural language processing (NLP) tasks [42] and then in computer vision [43]. In fact, the capacity of model to learn long dependencies is related to the length of the paths forward and backward signals. The shorter paths between any combination of positions in the input and output sequences allow to learn longer range dependencies[44]. For example, the work [42] compares the maximum path length between any two input and output positions in networks such as transformer, LSTM, and CNN and shows that transformer has the shortest paths and is capable of learning long range dependency.
Inspired by these facts, we propose in this paper a novel mix transformer for eye movement recognition, named EmMixformer. EmMixformer includes three modules, transformer, attention-LSTM, and Fourier-transformer. The eye movement signal is first split into fast and slow velocity signals, separately input to encoder for feature embedding. The embedding features are then input to the mix transformer module to learn a robust representation for classification. The transformer and attention LSTM aims at learning temporal dependencies within the eye movement signal, while the Fourier transformer learns a global feature representation. Our mix transformer can, therefore, learn complementary features. The contributions of our work are summarized as follows:
@@ II-A Siamese CNN block
The Siamese CNN block consists of two sub-networks, each stacking four one-dimensional convolutional modules for feature extraction. The size of the convolution kernel increases to obtain a progressively larger receptive field. Each module comprises a convolutional layer, a batch normalization layer, a ReLU layer, and an average pooling layer. The two subnets take respectively the preprocessed fast and slow eye movement signals as inputs and output the corresponding embedding feature representations. We then concatenate the output feature representations from the two subnets to obtain the complete feature representation. The $x$ and $y$ represent the horizontal and vertical coordinate sequences of eye movement data, and $n$ represents the sequence length.
@@ II-B Mix block
The Mix block comprises three modules, an attention LSTM (attLSTM), a Transformer, and a Fourier-transformer (Fourierformer). The three modules take the Siamese CNN block’s output as input to extract rich and comprehensive temporal-spatial features. Concretely, a temporal sequence from the Siamese CNN block is divided into a number of sub-sequences, which are input to attLSTM and transformer to extract the temporal features. The attLSTM module learns short-term dependencies within the temporal sequences, while the transformer module extracts long-term dependencies within the temporal sequences thanks to its self-attention mechanism [45]. Their combination, therefore, can learn a comprehensive temporal feature representation. The third module, Fourierformer, takes as input the temporal sequences to learn a global frequency-domain feature representation. As a result, our mix block extracts both a feature representation embedding temporal dependencies and a global feature representation, thus effectively improving the modeling of time-series eye movement classification.
@@
Transformer, introduced in 2017 by Vaswani et al [42], has become the state of the art for NLP [42] and more recently computer vision [45]. Unlike traditional sequence models such as recurrent neural networks (RNNs) or CNNs, Transformer relies solely on a key idea, the self-attention mechanism, eliminating the need for sequential processing. The self-attention mechanism allows the model to weigh the importance of different tokens in a sequence, and enables it to capture effectively long-range dependencies and context information. By attending to all tokens simultaneously, Transformer can process sentences in parallel, leading to significantly faster training and inference processing times, compared to existing sequential models [45].
@@
where $f_{t}$ , $i_{t}$ and $o_{t}$ are the forgetting gate, input gate and output gate, respectively. $\widetilde{C_{t}}$ , $C_{t}$ and $h_{t-1}$ are the candidate cell state, the cell state, and the hidden state at the last moment, respectively. $\odot$ denotes the element-by-element matrix multiplication and $\sigma$ represents the sigmoid activation function, respectively.
@@
$\mathcal{F}(X)\left(k\right)=\frac{1}{\sqrt{T}}\sum_{t=0}^{T-1}X(t)e^{-j\frac{2\pi}{T}kt}$ (24)
where $\mathcal{F}(X)\left(k\right)$ is the complex value of the $k$-th frequency component in the frequency domain, $X(t)$ is the complex value of the $t$th sampling point in the time domain, $T$ is the sequence length, and $\frac{1}{\sqrt{T}}$ ensures normalized consistency between the DFT and the inverse Fourier transform (IDFT).
Using the real part $Re(X)$ and the imaginary part $Im(X)$ of $\mathcal{F}(X)\left(k\right)$, we obtain the amplitude component $A(X)$ and the phase component $\varphi\left(X\right)$ by the following equations:
$A(X)\left(v\right)=\sqrt{Re\left(X\right)\left(k\right)^{2}+Im\left(X\right)\left(k\right)^{2}}$ (25)
$\varphi\left(X\right)\left(k\right)=arctan\left(\frac{Im\left(X\right)\left(k\right)}{Re\left(X\right)\left(k\right)}\right)$ (26)
Next, we extract the feature representation from amplitude component $A(X)\left(k\right)$ and phase component $\varphi(X)\left(k\right)$ using two independent transformers. The self attention operation is performed to capture the amplitude information $A_{transformer}$ and phase information $\varphi_{transformer}$, computed as:
@@
Subsequently, the feature representations of the amplitude and phase components are combined into complex numbers and transformed back into the spatial domain by IFFT
@@ III-A Dataset
CEM-{CJK*}UTF8gbsnⅠ[23] Tobii TX300 300 yes 22
CEM-{CJK*}UTF8gbsnⅡ[23] EyeLink 1000 1000 yes 32
@@ III-D Experimental results for short time interval
These experiments aim to assess the verification performance of our approach on short time interval data. In the three datasets, the participants provided data from two sessions at each round. As detailed in section III-C, we select data from the first session for training and the remaining data for testing. The EMglasses dataset comprises 203 sequences in the training set and testing set, the four sub-datasets (RAN, TEX, FXS, and HSS) in the the GazeBase dataset comprises 322 sequences in the training set and the testing set, while the JuDo1000 dataset comprises 450 sequences in the training set and 150 sequences in the testing set. Note that in all datasets, we split each sequence, collected for several seconds, into several sub-sequences for training and testing. The state of the art approaches, i.e. DEL[38], Expansion CNN[7], Dense LSTM[36] and DenseNet [40] are employed for benchmarking. The EER and FRR@FAR of each approach are listed in Table III, Table IV, Table VII, Table V, Table VI, and Table VIII. The corresponding ROC curves are shown in Fig. 7. It can be observed that our EmMixFormer outperforms the existing approaches and achieves the lowest verification error, i.e. 0.0673, 0.0635, 0.0801, 0.1578 on the four subsets of the GazeBase dataset, 0.0502 on the JuDo1000 dataset, and 0.1599 on our dataset. Also, we observe from Fig.7 that our proposed approach achieves the higher recognition accuracy w.r.t existing approaches at different FARs.
@@
EmMixformer 0.0673 0.0502 0.2032 0.4659
DEL[38] 0.1309 0.1680 0.6217 0.9087
Expansion CNN[7] 0.1437 0.1894 0.6267 0.9245
Dense LSTM[36] 0.1191 0.1365 0.5082 0.8382
DenseNet[40] 0.0839 0.0739 0.2691 0.5575
@@
EmMixformer 0.0635 0.0407 0.2603 0.6193
DEL[38] 0.1060 0.1128 0.5750 0.9141
Expansion CNN[7] 0.1362 0.1950 0.6977 1.0000
Dense LSTM[36] 0.0971 0.0945 0.4824 0.8456
DenseNet[40] 0.0736 0.0551 0.3293 0.7175
@@
EmMixformer 0.0801 0.0680 0.2818 0.6087
DEL[38] 0.1436 0.2066 0.7383 0.9645
Expansion CNN[7] 0.1500 0.2340 0.7277 0.9902
Dense LSTM[36] 0.1161 0.1329 0.5529 0.8846
DenseNet[40] 0.0885 0.0807 0.3513 0.7045
@@
EmMixformer 0.1578 0.2549 0.7654 0.9539
DEL[38] 0.2229 0.5022 0.9358 0.9936
Expansion CNN[7] 0.1950 0.4051 0.9067 0.9903
Dense LSTM[36] 0.2191 0.4047 0.8831 0.9832
DenseNet[40] 0.1666 0.2979 0.8323 0.9821
@@
EmMixformer 0.1599 0.2572 0.7888 0.9711
DEL[38] 0.1853 0.3565 0.8877 0.9900
Expansion CNN[7] 0.2140 0.4013 0.8654 0.9821
Dense LSTM[36] 0.2069 0.4545 0.8928 1.0000
DenseNet[40] 0.1892 0.3444 0.8131 0.9812
@@ III-E Experimental results for long time intervals
TABLE IX: Results of long-term interval experiments on the RAN database
@@
EmMixformer 0.0801 0.1428 0.1803 0.1793 0.1853 0.2279
DEL[38] 0.1436 0.2215 0.2594 0.2926 0.2459 0.2907
Expansion CNN[7] 0.1500 0.2328 0.2769 0.2395 0.2812 0.3057
Dense LSTM[36] 0.1161 0.1939 0.2535 0.2791 0.2322 0.2840
DenseNet[40] 0.0885 0.1775 0.2161 0.2032 0.2047 0.2592
@@ III-F Ablation experiment
Our our EmMixFormer consists of siamese CNN and Mix block, where Mix block includes three modules: Attention LSTM, Transformer, and Fourier Transformer (as shown in Fig.2). To investigate the effect of each module w.r.t recognition accuracy improvement, we conducted ablation experiments on the round 1 data from the RAN sub-dataset in GazeBase. To ease the description, we remove Mix block from EmMixFormer to obtain the basic siamese CNN. Then, we gradually add attention LSTM, Transformer, and Fourier Transformer, with the resulting models denoted as Siamese CNN+Transformer, Siamese CNN+LSTM+Transformer and Siamese CNN+attLSTM+Transformer, respectively. Afterwards, the Fourier Transformer, is included to the attLSTM+Transformer, with the resulting model represented as Siamese CNN+attLSTM+Transformer+Fourierformer, i.e. EmMixFormer. The verification error rates of these ablation schemes are shown in Table X and Fig.8. The experimental results demonstrate that incorporating Transformer into siamese CNN increases the accuracy, which implies that Transformer can learn dependency features to improve recognition performance. Also, we can observe that the combination of the Attention LSTM and the Fourier Transformer allows to improve the identification accuracy of simple Transformer. Besides, comparing Transformer with LSTM+Transformer, we observe that the LSTM can provide compensation features to improve the performance of Transformer, which is supported by the work [56]. From Table X and Fig.8, we also observe that attLSTM+Transformer achieves lower EER than LSTM+Transformer, which implies that the attention mechanism improves LSTM performance. Finally, the performance becomes optimal when attLSTM+Transformer is combined with Fourier Transformer, which demonstrates that the attention mechanism in the frequency domain captures global dependencies for feature representation.
TABLE X: Results of the ablation experiments on the RAN database
Method EER FRR@FAR
$10^{-1}$ $10^{-2}$ $10^{-3}$
Siamese CNN 0.1594 0.2498 0.7091 0.9900
+Transformer 0.1214 0.1491 0.6167 0.9133
LSTM+Transformer 0.0880 0.0776 0.4148 0.8204
attLSTM+Transformer 0.0936 0.0887 0.3720 0.7377
EmMixformer 0.0801 0.0680 0.2818 0.6087